Reports 1-1 of 1 Clear search Modify search
fabian.arellano - 13:52 Thursday 18 November 2021 (18933) Print this report
Comment to In-vacuum cable health check at outer side of flanges. (18191)

GAS  filter  LVDT  card  replacement

Summary:  the  new  LVDT  card  seems  to  be  working  fine,  but  the  calibration  factors  and  offsets  seem  to  require  adjustment.

After the  replacement  of  the  LVDT  card  for  the  GAS  filters,  I  chracterized  the  health  of  the  LVDTs  using  transfer  fuction  measurements.

Bottom  Filter

The position  of  the  keystone  given  by the  LVDT  is  close  to  what I  was  expecting.  Before  the  replacement  of  the  card  it  was  -186  um  and  after the  replacement  it was -179  um,  as  shown  in  the  first  screeshot.  The  second  screenshot shows  the  transfer  function.  Given the  logaritimic  scale,  the  amplitude seems  very  similar  to  what  it  was  before,  namely,  (  reference  /  measurement ) = [ 0.0358 (um/cnt) /  0.0228 (um/cnt) ]  =  1.570  (below  the  lowest  resonance  frequency).

Then  I moved the keystone close to its setpoint, checked the vertical OSEMs  were  within  range ( IM-V  =  136  um)  and  measured  the  IM-V  transfer  function. The  result  was what  I  was  expecting.  See  the  third  screenshot.

Standard  Filter  F1

The  position  of  the keystone  given  by  the LVDT  changed  a  lot with  respect  to  what  it  was  before.  It  went  from  1007  um  to  2387  um  (first  screenshot  also).  The transfer  function,  in the  fouth  screenshot,  looks fine,  although  larger than the  reference.  We  should check  the  last transfer  function  measured  before  the  LVDT  began  having  problems  to  compare  to the  current  one,  but  for  now  let's  use  the  reference:   (  reference  /  measurement ) = [ 0.0340 (um/cnt) /  0.0490 (um/cnt) ]  =  0.710 .

Top  Filter  F0

As  in t he  case  of  F1,  the  position  of  the keystone  given  by  the LVDT  changed  a  lot with  respect  to  what  it  was  before.  It  went  from  -764 um  to -2479  um  (first  screenshot  also).  The transfer  function,  on  the left-hand  side  of  the  fifth  screenshot,  looks fine.  The  amplitude is  larger than  the  reference,  nevetheless,  a  previous  transfer  function,  measured  on  the  15th  of  April  2019  and  shown  on  the  right-hand  side,  also  shows  a  larger  amplitude than  the  reference.  I'll use  this  measurement  for  comparison  instead  of  the  reference:  (  ref_april2019  /  measurement ) = [ 0.0585 (um/cnt) /  0.0775 (um/cnt) ]  =  0.754 .


Although  the  transfer  functions  are  fine  in  terms  of  shape,  it  seems  that the  calibration  factors  (in  units  of  um/count)  and  offsets (in  units  of  counts)   need  additional  adjustment.  In  principle,  because  we  adjusted  the  resistances  in the  LVDT  card,  this  should  not be  necessary.  However,  practice  says  otherwise. We might  have  made  a  mistake in  adjusting  the  resistances.  Another  element  to  consider  that the  the  new  boards  are  slightly  different  and  I  don't  know  whether  this  would  affect the gains  and  phases  that  determine  the  offset  and  calibration  factor.

Images attached to this comment
Search Help